# Antonio Raiola

### v1

@@ 4.1 Panels with $N\to\infty$ and $T$-fixed
When $T$ is fixed and $N$ grows, the only available slope homogeneity test among those considered is the LM-type test developed by BRS. The restrictions on the regressors can be relaxed, as in Assumption 3 in the Appendix, which also contains the corresponding version of Theorem 1. In this scenario, the boundary between trivial and nontrivial power is governed by the rate
@@
When the size of the alternative group is proportional to the total cross-section, $m_{0}>0$ (or $M\!\asymp\!N$), the rate in (6) reduces to $\gamma\!\asymp\!N^{1/4}$, which coincides with the local-power behavior established in BRS. More interestingly, the rate in (6) marks the boundary at which the test begins to exhibit power. As long as $M$ diverges at least at the rate $N^{1/2}$, the $LM$ test retains nontrivial local power even when $M/N\!\to\!0$. Essentially, this result identifies the minimal detectable group size as being of order $N^{1/2}$: below this threshold, grouped heterogeneity is undetectable.
@@ 4.2 Panels with increasing number of groups
The main message of our analysis will remain valid even if we allow the number of groups to grow with $N$. Consider
$H_{2,n}:\beta_{g(i)}=\begin{cases}\beta,&\mbox{if }i\in G_{1},\\ \beta+\lambda_{p}/\gamma,&\mbox{if }i\in G_{p}\mbox{ and }p=2,...,P,\end{cases}$
@@
The analysis of local power when $P\to\infty$ as $N\to\infty$ is complicated by the fact that the number of individuals in each group $\{G_{2},...,G_{P}\}$ can also be a function of $N$. Hence, the combination of both number of groups and number of individuals in each group plays a key role to obtain the expansion rates under which the test statistics have nontrivial power.
@@
$\lim_{N,T\to\infty}\frac{1}{\sqrt{2K}}\left(\sum_{p=2}^{P}c_{p}^{2}\lambda_{p}^{\prime}\left(Q^{(G_{p})}\lambda_{p}\right)-\sum_{p,q=2}^{P}m_{p}c_{q}^{2}\lambda_{p}\left(Q^{(G_{p})}Q^{-1}Q^{(G_{q})}\right)\lambda_{q}\right),$
@@
respectively, where for all p=2,,Pp=2,...,P, $M_{p}/N\to m_{p}\in[0,1)$ and $M_{p}T/(\gamma\sqrt{N})\to c_{p}^{2}\in[0,c^{2}]$, with $c=\mathbb{M}T/(\sqrt{N}\gamma^{2})$.
Corollary 1 generalize the group structure in Theorem 1. To see that Corollary 1 nests previous results assume that $P$ is fixed and $M^{\ast}=\max_{2\leq p\leq P}M_{p}\to\infty$ as $N\to\infty$. Then by replacing $M=M^{\ast}$ the results in Theorem 1 follows. Further, this corollary also nests the case when $P\to\infty$ as $N,T\to\infty$ and the number of individuals in groups $\{G_{2},...,G_{P}\}$ is fixed and bounded (i.e., $\sup_{2\leq p\leq P}M_{p}=\kappa<\infty$). Then, the expansion rates is given by $PT/(\sqrt{N}\gamma^{2})\to c>0$ with $N,T\to\infty$. Finally, in the general case when both $P\to\infty$ and $M_{p}\to\infty$ for at least one p=2,,Pp=2,...,P, the corresponding expansion rate will be driven by the overall “volume” of individual in $\{G_{2},...,G_{P}\}$, defined by $\mathbb{M}$.
@@ 5 Monte Carlo Study
We evaluate the finite-sample performance of the tests in terms of their power, relating to and verifying the theoretical results provided in the main text. We consider the three test statistics ($\Delta$, $LM$, and $\mathcal{J}$) under a data-generating process adapted from PY to fit the context of grouped patterns; i.e., for each $i=1,\ldots,N$ and $t=1,\ldots,T$:
$y_{it}$ $=x_{it}\beta_{i}+\alpha_{i}+\epsilon_{it},$
$x_{it}$ $=\rho_{i}x_{it-1}+\sqrt{1-\rho_{i}^{2}}\,v_{it},$
where $\epsilon_{it}\overset{iid}{\sim}\mathcal{N}(0,\sigma_{i}^{2})$, $v_{it}\overset{iid}{\sim}\mathcal{N}(0,\sigma_{ix}^{2})$, $\alpha_{i}\overset{iid}{\sim}\mathcal{N}(1,1)$, and $\rho_{i}\overset{iid}{\sim}\mathcal{U}(0.05,0.95)$, with $\sigma_{i}^{2}\sim\mathcal{X}^{2}(2)/2$ and $\sigma_{ix}^{2}\sim\mathcal{X}^{2}(1)$. The first 50 observations of $x_{it}$ ($t=-49,\ldots,0$) are discarded to mitigate the effect of initial values. Group-specific slopes are generated as:
$\beta_{i}=\begin{cases}1,&\text{if }i\in G_{1},\\ 1+\lambda(1+e_{p}),&\text{if }i\in G_{p},\end{cases}$
where $p=2,\ldots,P$ and $e_{p}\overset{iid}{\sim}\mathcal{U}(-h,h)$. The first group, $G_{1}$, contains $N-\sum_{p=2}^{P}M_{p}$ individuals, while the remaining P1P-1 groups $\{G_{2},\ldots,G_{P}\}$ represent alternative groups of size $M_{p}$ each. The parameters $\lambda$ and $h$ control the heterogeneity between the dominant and remaining groups, and across alternative groups, respectively.
We consider two simulation designs to examine how the cardinality of the alternative group and the magnitude of slope differences affect statistical power. Each design considers $(N,T)=(100,10)$ and $(100,100)$, corresponding to panels of small and large time dimensions, respectively:
@@
    Two-group design: $P=2$, $h=0$, varying the size of the alternative group $M_{2}$ for $\lambda\in\{0.1,0.3,0.5\}$.
@@
    Multiple-group design: $P\in\{3,5,10\}$, where alternative groups size are randomly assigned subject to $\sum_{p=2}^{P}M_{p}=\mathbb{M}$, with varying $\mathbb{M}$. Here, $\lambda=0.1$ and $h\in\{0.2\}$, focusing on the effect of the overall mass of alternative groups $\mathbb{M}$.
The data-generating process includes individual constants, removed through appropriate demeaning procedures for each test statistic. Specifically, for the $LM$ test, we use orthogonal deviations as in Arellano and Bover (1995), following Remark 5.1 in BRS. For the $\Delta$ and $\mathcal{J}$ tests, we apply the classical demeaning procedure, pre-multiplying $y_{i}$ and $x_{i}$ by $M_{0}=I_{T}-T^{-1}\textbf{i}_{T}\textbf{i}_{T}^{\prime}$, where $\textbf{i}_{T}$ is a $T\times 1$ vector of ones, as suggested in PY and Lu and Su (2017). All simulations report rejection rates at the 5% significance level, based on 1,000 Monte Carlo replications. The results are presented in Figures 1 and 2, respectively.
Two Groups $N=100$, $T=10$
Two Groups $N=100$, $T=100$
@@
The two-group scenario results (Figure 1) closely mirror our theoretical findings in Theorem 1. The rejection rates increase monotonically with the alternative group size $M_{2}$ and with the slope deviation parameter $\lambda$. All tests tend to have a power close to trivial when $M_{2}$ is smaller than $\sqrt{N}$, but the power increases significantly after this threshold and when $T$ grows.
In the multiple-group experiment (Figure 2), the decisive role is played by the number of individuals in the alternative groups, $\mathbb{M}$. Tests exhibit substantial power whenever $\mathbb{M}$ grows at least proportionally to $\sqrt{N}$, consistent with Corollary 1. Notice that the tests power decreases with larger $P$, but this has only a secondary impact once $\mathbb{M}$ and/or $T$ is large enough.
$P$ Groups $N=100$, $T=10$
$P$ Groups $N=100$, $T=100$
Figure 2: Power curves of slope homogeneity tests with $P$ groups.
@@ 6 Conclusions
We propose repurposing existing slope-homogeneity tests to detect grouped patterns in panels. The power of the tests is evaluated under a sequence of doubly local alternatives, linking the degree of heterogeneity across groups and each group size to its impact on test performance. Our analysis shows that slope-homogeneity tests are effective in detecting such structures, including sparse groups whose cardinality is negligible relative to the cross-sectional size. Monte Carlo simulations support similar conclusions in finite samples.
@@ Appendix A Proofs
We derive the asymptotic distribution of the three type of tests in Theorems 1 under $H_{1,n}$, in the large panel setting, where $N,T\to\infty$. The rates stated in Corollaries 1 and 2 (at the end of the Appendix) follow immediately by setting $M=K_{N}$ and $T=1$, respectively. Additional remarks on the derivation of the noncentrality parameters in Corollary 1 are provided at the end of this section.
@@
$\mathop{\rm plim}_{N,T\to\infty}\frac{1}{\sqrt{N}\gamma^{2}}\sum_{p=2}^{P}\lambda_{p}^{\prime}\left(\sum_{i\in G_{p}}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\lambda_{p}=\lim_{N,T\to\infty}\sum_{p=2}^{P}\lambda_{p}^{\prime}SQ^{(G_{p})}\lambda_{p}\left(\frac{M_{p}T}{\sqrt{N}\gamma^{2}}\right)+o_{p}(1).$
@@
$\frac{1}{\sqrt{N}}R_{NT}$ $=-\frac{1}{\sqrt{N}\gamma^{2}}\left(\sum_{p=2}^{P}\lambda_{p}^{\prime}\left(\sum_{i\in G_{p}}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\right)\frac{S}{NT}\left(\sum_{q=2}^{P}\lambda_{p}\left(\sum_{i\in G_{q}}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\right)+o_{p}(1)$
$=\sum_{p=2}^{P}\sum_{q=2}^{P}\lambda_{p}^{\prime}S^{(G_{p})}SS^{(G_{q})}\lambda_{q}\left(\frac{M_{p}}{N}\right)\left(\frac{M_{p}T}{\sqrt{N}\gamma^{2}}\right)+o_{p}(1)$

### v2

@@ 4.1 Panels with $N\to\infty$ and $T$-fixed
When $T$ is fixed and $N$ grows, the LM-type test of BRS is the only one available. The restrictions on the regressors can be relaxed, as in Assumption 3 in the Appendix, which also contains the corresponding version of Theorem 1. In this scenario, the boundary between trivial and nontrivial power is governed by the rate
@@
When the alternative group is proportional to the total cross-section ($m_{0}>0$, or $M\!\asymp\!N$), the rate in (6) reduces to $\gamma\!\asymp\!N^{1/4}$, matching the local-power behavior established in BRS. More interestingly, this rate marks the boundary at which the test begins to exhibit power. If $M$ diverges at least as fast as $N^{1/2}$, the $LM$ test retains nontrivial local power even when $M/N\!\to\!0$. Essentially, the minimal detectable group size is of order $N^{1/2}$: below this threshold, grouped heterogeneity is undetectable.
@@ 4.2 Panels with increasing number of groups
The main message of our analysis remain valid even if the number of groups is allowed to grow with $N$. Consider
$H_{2,n}:\beta_{g(i)}=\begin{cases}\beta,&\mbox{if }g(i)=G_{1},\\ \beta+\lambda_{p}/\gamma,&\mbox{if }g(i)=G_{p}\mbox{ and }p=2,...,P,\end{cases}$
@@
Analyzing the local power when $P\to\infty$ as $N\to\infty$ is complicated because the size of each group $\{G_{2},\dots,G_{P}\}$ can also depend on $N$. Both the number of groups and the number of individuals per group determine the expansion rates under which the test statistics exhibit nontrivial power.
@@
$\lim_{N,T\to\infty}\frac{1}{\sqrt{2K}}\left(\sum_{p=2}^{P}c_{p}^{2}\lambda_{p}^{\prime}Q^{(G_{p})}\lambda_{p}-\sum_{p,q=2}^{P}m_{p}c_{q}^{2}\lambda_{p}\left(Q^{(G_{p})}Q^{-1}Q^{(G_{q})}\right)\lambda_{q}\right),$
@@
respectively, where for all p=2,,Pp=2,...,P, $M_{p}/N\to m_{p}\in[0,1)$ and $M_{p}T/(\gamma^{2}\sqrt{N})\to c_{p}^{2}\in[0,c^{2}]$, with $c=\mathbb{M}T/(\sqrt{N}\gamma^{2})$.
Corollary 1 generalizes Theorem 1. If $P$ is fixed and $M^{\ast}=\max_{2\leq p\leq P}M_{p}\to\infty$ as $N\to\infty$, replacing $M$ with $M^{\ast}$ yields Theorem 1. If $P\to\infty$ while $\sup_{2\leq p\leq P}M_{p}=\kappa<\infty$, the expansion rate satisfies $PT/(\sqrt{N}\gamma^{2})\to c>0$ as $N,T\to\infty$. More generally, when both $P\to\infty$ and $M_{p}\to\infty$ for at least one $p$, the expansion rate is determined by the total “volume” of individuals in $\{G_{2},\dots,G_{P}\}$, denoted as $\mathbb{M}$.
@@ 5 Monte Carlo Study
We evaluate the finite-sample performance of the tests by comparing their power and assessing the theoretical results presented in the main text. We consider the three test statistics ($\Delta$, $\mathcal{J}$, and $LM$) under a data-generating process adapted from PY to accommodate grouped slope patterns; i.e., for each $i=1,\ldots,N$ and $t=1,\ldots,T$:
$y_{it}$ $=x_{it}^{\prime}\beta_{g(i)}+\alpha_{i}+\epsilon_{it},$
$x_{it,k}$ $=\rho_{i,k}x_{it-1,k}+\sqrt{1-\rho_{i,k}^{2}}\,v_{it,k},\qquad k=1,\ldots,K,$
where $x_{it}=(x_{it,1},\ldots,x_{it,K})^{\prime}$, $\epsilon_{it}\overset{iid}{\sim}\mathcal{N}(0,1)$, $v_{it}\overset{iid}{\sim}\mathcal{N}(0,\sigma_{ix}^{2})$, $\alpha_{i}\overset{iid}{\sim}\mathcal{N}(1,1)$, and $\rho_{i}\overset{iid}{\sim}\mathcal{U}(0.05,0.95)$, with $\sigma_{ix}^{2}\sim\chi^{2}_{1}$. The first 50 observations of $x_{it}$ ($t=-49,\ldots,0$) are discarded to mitigate the effect of initial values. Group-specific slopes, $\beta_{g(i)}=(\beta_{g(i),1},...,\beta_{g(i),K})^{\prime}$, are generated as:
$\beta_{g(i),k}=\begin{cases}1,&\text{if }g(i)=G_{1},\\ 1+\lambda(1+e_{p}),&\text{if }g(i)=G_{p},\end{cases}$
for $k=1,\ldots,K$, where $p=2,\ldots,P$ and $e_{p}\overset{iid}{\sim}\mathcal{U}(-h,h)$. The dominant group $G_{1}$ contains $N-\sum_{p=2}^{P}M_{p}$ individuals, while the remaining P1P-1 groups $\{G_{2},\ldots,G_{P}\}$ represent alternative groups of size $M_{p}$. The parameters $\lambda$ and $h$ control, respectively, the degree of heterogeneity relative to the dominant group and the dispersion across alternative groups.
We consider two simulation designs to examine how the size of the alternative group and the magnitude of slope differences affect power:
@@
    Two-group design: $P=2$, $h=0$, varying the size of the alternative group $M_{2}$ for $\lambda\in\{0.1,0.2,0.3\}$.
@@
    Multiple-group design: $P\in\{3,5,10\}$, where alternative groups size are randomly assigned subject to $\sum_{p=2}^{P}M_{p}=\mathbb{M}$, with varying $\mathbb{M}$. Here, $\lambda=0.2$ and $h=0.2$, focusing on the effect of the overall mass of alternative groups $\mathbb{M}$.
Each design is run at $(N,T)=(100,10)$ and $(100,100)$, corresponding to short and long panels, and at $K=1$ and $K=4$, corresponding to single and multiple regression.
In both designs, the $\mathcal{J}$ test is severely biased when $T$ is small, especially for $K=4$, and its rejection rate far exceeds that of the other two tests. To compare the three tests on equal footing, we use size-adjusted critical values obtained by simulating the null distribution of each test with $R_{2}=10000$ replications.
The data-generating process includes individual effects, which are removed through appropriate demeaning procedures for each test statistic. For the $LM$ test, we employ orthogonal deviations as in Arellano and Bover (1995), following Remark 5.1 in BRS. For the $\Delta$ and $\mathcal{J}$ tests, we apply the standard within transformation, pre-multiplying $y_{i}$ and $x_{i}$ by $M_{0}=I_{T}-T^{-1}\mathbf{i}_{T}\mathbf{i}_{T}^{\prime}$, where $\mathbf{i}_{T}$ is a $T\times 1$ vector of ones, as suggested in PY and Lu and Su (2017).
Two Groups: $K=1$, $N=100$, $T=10$
Two Groups: $K=1$, $N=100$, $T=100$
Two Groups: $K=4$, $N=100$, $T=10$
Two Groups: $K=4$, $N=100$, $T=100$
@@
$P$ Groups: $K=1$, $N=100$, $T=10$
$P$ Groups: $K=1$, $N=100$, $T=100$
$P$ Groups: $K=4$, $N=100$, $T=10$
$P$ Groups: $K=4$, $N=100$, $T=100$
Figure 2: Power curves of slope homogeneity tests with two groups.
All simulations report rejection frequencies at the 5% significance level based on 1,000 Monte Carlo replications. Results for the two-group design are shown in Figure 1, and those for the multiple-group design in Figure 2.
The two-group results closely mirror our theoretical findings in Theorem 1. Rejection rates increase monotonically with the alternative group size $M$ and with the slope deviation parameter $\lambda$. All tests tend to have a power close to trivial when $M$ is smaller than $\sqrt{N}$, but the power increases significantly after this threshold and as $T$ grows.
In the multiple-group experiment, the decisive role is played by the number of individuals in the alternative groups, $\mathbb{M}$. Tests exhibit substantial power when $\mathbb{M}$ grows at least proportionally to $\sqrt{N}$, consistent with Corollary 1. Power declines slightly as $P$ increases, due to greater dispersion in group-specific slopes, but this effect is secondary once $\mathbb{M}$ or $T$ is sufficiently large.
In most cases, all three tests show comparable power. The $J$ and $LM$ tests, however, tend to gain power somewhat more rapidly than the $\Delta$ test, especially in the short-panel designs with $T=10$ and $K=4$. This divergence is much less visible once $T$ increases. When $T=100$, the three procedures behave in a way that is broadly consistent with the asymptotic comparison in Section 4.
@@ 6 Conclusions
We derive the detection boundary of standard slope-homogeneity tests under small-group heterogeneity. The power of the tests is evaluated under a sequence of doubly local alternatives, linking the degree of heterogeneity across groups and each group size to its impact on test performance. Our analysis shows that slope-homogeneity tests are effective in detecting such structures, including sparse groups whose cardinality is negligible relative to the cross-sectional size. Monte Carlo simulations support similar conclusions in finite samples.
@@ Appendix A Proofs
We derive the asymptotic distribution of the three type of tests in Theorems 1 under $H_{1,n}$, in the large panel setting, where $N,T\to\infty$. The rates stated in Corollaries 1 and 2 (at the end of the Appendix) follow immediately by setting $M=\mathbb{M}$ and $T=1$, respectively. Additional remarks on the derivation of the noncentrality parameters in Corollary 1 are provided at the end of this section.
@@
$\mathop{\rm plim}_{N,T\to\infty}\frac{1}{\sqrt{N}\gamma^{2}}\sum_{p=2}^{P}\lambda_{p}^{\prime}\left(\sum_{i\in G_{p}}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\lambda_{p}=\lim_{N,T\to\infty}\sum_{p=2}^{P}\lambda_{p}^{\prime}Q^{(G_{p})}\lambda_{p}\left(\frac{M_{p}T}{\sqrt{N}\gamma^{2}}\right)+o_{p}(1).$
@@
$\frac{1}{\sqrt{N}}R_{NT}$ $=-\frac{1}{\sqrt{N}\gamma^{2}}\left(\sum_{p=2}^{P}\lambda_{p}^{\prime}\left(\sum_{i\in G_{p}}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\right)\frac{Q^{-1}}{NT}\left(\sum_{q=2}^{P}\lambda_{p}\left(\sum_{i\in G_{q}}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\right)+o_{p}(1)$
$=\sum_{p=2}^{P}\sum_{q=2}^{P}\lambda_{q}^{\prime}Q^{(G_{p})}Q^{-1}Q^{(G_{q})}\lambda_{q}\left(\frac{M_{p}}{N}\right)\left(\frac{M_{p}T}{\sqrt{N}\gamma^{2}}\right)+o_{p}(1)$